# Gaussian Process-Based Prediction and Control of Hammerstein-Wiener Systems

### v1

@@ II-A Data-Driven Prediction as Linear Systems
$\left[\Gamma_{1}\ \Gamma_{2}\right]:=H_{\bar{y}\text{f}}\mathrm{col}\left(H_{\bar{u}},H_{\bar{y}\text{p}}\right)^{\dagger}\in\mathbb{R}^{n_{y}L^{\prime}\times(n_{u}L+n_{y}L_{0})}.$ (3)
Proof:
This is a direct extension of the data-driven simulation result, originally proposed in [13, Proposition 1]. ∎
Remark 2
Equation (2) can be seen as a nonlinear multi-step-ahead ARX model:
$\mathbf{y}_{\text{f}}=\Phi^{-1}\left(\left[\Gamma_{1}\ \Gamma_{2}\right]\mathrm{col}\left(\Psi(\mathbf{u}),\Phi(\mathbf{y}_{\text{p}})\right)+\left[-\Gamma_{2}\ \mathbb{I}\right]\mathbf{e}\right).$
For linear systems, this predictor is known as the subspace predictor [27], and is equivalent to the solution to the least-squares problem [28]:
$\underset{\Gamma_{1},\Gamma_{2}}{\text{min}}\ \left\lVert H_{\bar{y}\text{f}}-\left[\Gamma_{1}\ \Gamma_{2}\right]\mathrm{col}\left(H_{\bar{u}},H_{\bar{y}\text{p}}\right)\right\rVert_{F}^{2}.$
Remark 3
A strategy to analyze Hammerstein-Wiener systems with unknown input and output nonlinearities is to assume that finite-dimensional basis functions are available for the nonlinearities [19], i.e., $\psi(\cdot)=:\sum_{i=1}^{r}a_{i}\psi_{i}(\cdot)=:\left(\mathbf{a}^{\top}\otimes\mathbb{I}\right)\boldsymbol{\psi}(\cdot)$, $\phi(\cdot)=:\sum_{i=1}^{q}b_{i}\phi_{i}(\cdot)=:\left(\mathbf{b}^{\top}\otimes\mathbb{I}\right)\boldsymbol{\phi}(\cdot)$. This leads to
@@
where $\bar{B}:=B\left(\mathbf{a}^{\top}\otimes\mathbb{I}\right)$ and $\bar{D}:=D\left(\mathbf{a}^{\top}\otimes\mathbb{I}\right)$. By considering augmented inputs $\boldsymbol{\psi}(u_{k})$ and outputs $\boldsymbol{\phi}(y_{k})$, (4) admits linearity in the sense that for any two input-output trajectories of (4), their linear combinations are still valid trajectories. However, as also pointed out in [19], (4) is not a standard linear model since the operator $\left(\mathbf{b}^{\top}\otimes\mathbb{I}\right)\in\mathbb{R}^{n_{y}\times n_{y}q}$ is not invertible. This prohibits the use of Willems’ fundamental lemma to quantify all possible input-output trajectories and provide data-driven predictions when unknown nonlinear output uncertainties exist. In addition, the assumption of having finite-dimensional basis function knowledge is restrictive and impractical for many applications.
III Gaussian Process Data-Driven Prediction
III-A Gaussian Process Regression
Since the input and output nonlinearities are unknown, they are modeled as Gaussian processes with prior distributions:
Remark 4
Most standard kernel structures include a scaling parameter as a hyperparameter. In our setting, as discussed in Remark 1, this scaling parameter can be fixed to 1 without loss of generality. Similarly, the mean functions $m_{u}$ and $m_{y}$ can also be normalized.
We are interested in jointly learning the linear mapping $\Gamma_{1},\Gamma_{2}$ and the nonlinearities $\psi(\cdot)$ and $\phi(\cdot)$. Although the prior distribution of $\Gamma_{1},\Gamma_{2}$ can be derived from (3) and the Gaussian prior (5), it is dependent on $\psi(\cdot)$ and $\phi(\cdot)$ in a nonlinear way, and thus the Gaussianity is not preserved. Instead, we opt for directly learning $\Gamma_{1},\Gamma_{2}$ in a hierarchical model as will be detailed later.
Remark 5
@@ III-A Gaussian Process Regression
We first briefly summarize the standard procedure of Gaussian process regression [29]. Consider the problem of learning an unknown function $f:\mathbb{R}^{n_{\eta}}\rightarrow\mathbb{R}^{n_{\chi}}$ from data $\left(\boldsymbol{\boldsymbol{\eta}}^{d},\boldsymbol{\chi}^{d}\right)=\left(\eta_{k}^{d},\chi_{k}^{d}\right)_{k=1}^{M}$, where $\chi^{d}_{k}=f(\eta^{d}_{k})+\epsilon_{k}$ and $\epsilon_{k}$ is zero-mean i.i.d. Gaussian with covariance $\Sigma$. The function $f$ is modeled as a Gaussian process with prior mean $m(\cdot)$, prior covariance $k(\cdot,\cdot;\Theta)$, and hyperparameters $\Theta$, i.e.,
$f(\cdot)\sim\mathcal{GP}\left(m(\cdot),k(\cdot,\cdot;\Theta)\right).$
To predict $f(\eta)$ at any query point $\eta$, the joint distribution of $f(\eta)$ and $\boldsymbol{\chi}^{d}$ is given by the following Gaussian distribution:
$\begin{bmatrix}f(\eta)\\ \boldsymbol{\chi}^{d}\end{bmatrix}\sim\mathcal{N}\left(\begin{bmatrix}m(\eta)\\ \mathbf{m}(\boldsymbol{\eta}^{d})\end{bmatrix},\begin{bmatrix}k(\eta,\eta)&\mathbf{k}^{\top}(\boldsymbol{\eta}^{d},\eta)\\ \mathbf{k}(\boldsymbol{\eta}^{d},\eta)&K(\boldsymbol{\eta}^{d},\boldsymbol{\eta}^{d})+\bar{\Sigma}\end{bmatrix}\right),$
@@
$f(\eta)|\boldsymbol{\chi}^{d}\sim\mathcal{N}\left(m_{\text{p}}(\eta),k_{\text{p}}(\eta)\right),$ (9)
@@
$m_{\text{p}}(\eta)$ $:=m(\eta)+\mathbf{k}^{\top}(\boldsymbol{\eta}^{d},\eta)\left(K+\bar{\Sigma}\right)^{-1}\left(\boldsymbol{\chi}^{d}-\mathbf{m}(\boldsymbol{\eta}^{d})\right),$ (10)
$k_{\text{p}}(\eta)$ $:=k(\eta,\eta)-\mathbf{k}^{\top}(\boldsymbol{\eta}^{d},\eta)\left(K+\bar{\Sigma}\right)^{-1}\mathbf{k}(\boldsymbol{\eta}^{d},\eta),$ (11)
where the dependence of $K$ is omitted for brevity. Then, $f(\eta)$ is predicted as the posterior mean $m_{\text{p}}(\eta)$.
@@
$\Theta$ $=\text{arg}\underset{\Theta}{\text{max}}\ p\left(\boldsymbol{\chi}^{d}|\boldsymbol{\eta}^{d},\Theta\right)$ (12)
$=\text{arg}\underset{\Theta}{\text{min}}\ \mathrm{logdet}\left(\Lambda(\Theta)\right)+\delta^{\top}\Lambda^{-1}(\Theta)\delta,$ (13)
where $\delta:=\boldsymbol{\chi}^{d}-\mathbf{m}(\boldsymbol{\eta}^{d})$ and
$\Lambda(\Theta):=K(\boldsymbol{\eta}^{d},\boldsymbol{\eta}^{d};\Theta)+\bar{\Sigma}.$ (14)
III-B Kernel Design and Posterior Prediction
For this particular data-driven prediction problem, consider the trajectory characterization equation (2). Let $\eta=\mathrm{col}\left(\mathbf{u},\mathbf{y}\right)$ and
$f(\eta)=\left[\Gamma_{1}\ \Gamma_{2}\ -\mathbb{I}\right]\mathrm{col}\left(\Psi(\mathbf{u}),\Phi(\mathbf{y}_{\text{p}}),\Phi(\mathbf{y}_{\text{f}})\right),$ (15)
where $n_{\eta}=(n_{u}+n_{y})L$ and $n_{\chi}=n_{y}L^{\prime}$. We would like to train a Gaussian process model of $f(\eta)$ by using data $\left(\boldsymbol{\boldsymbol{\eta}}^{d},\boldsymbol{\chi}^{d}\right)$ where $M=N-L+1$,
@@ III-B Kernel Design and Posterior Prediction
$\bar{\Sigma}$ $=\mathbb{I}\otimes\mathrm{cov}\left(\left[-\Gamma_{2}\ \mathbb{I}\right]\mathbf{e}\right)=\sigma^{2}\left(\mathbb{I}\otimes\left(\Gamma_{2}\Gamma_{2}^{\top}+\mathbb{I}\right)\right).$ (17)
This is known as an implicit Gaussian process model [26].
The prior mean and covariance functions $m(\cdot)$ and $k(\cdot,\cdot)$ are then derived from the Gaussian process characterization of $\psi(\cdot)$ and $\phi(\cdot)$ in (5).
Consider two input-output trajectories $\eta^{\prime}=\mathrm{col}\left(\mathbf{u}^{\prime},\mathbf{y}^{\prime}\right)$ and $\eta^{\prime\prime}=\mathrm{col}\left(\mathbf{u}^{\prime\prime},\mathbf{y}^{\prime\prime}\right)$. The mean function is given by
$m(\eta^{\prime})=\left[\Gamma_{1}\ \Gamma_{2}\ -\mathbb{I}\right]\mathrm{col}\left(\mathbf{m}_{u}(\mathbf{u}^{\prime}),\mathbf{m}_{y}(\mathbf{y}^{\prime})\right),$ (18)
where $\mathbf{m}_{u}(\cdot)$, $\mathbf{m}_{y}(\cdot)$ apply $m_{u}(\cdot)$, $m_{y}(\cdot)$ elementwise, respectively. The covariance matrix between $\Psi(\mathbf{u}^{\prime})$ and $\Psi(\mathbf{u}^{\prime\prime})$ is given by $\mathrm{cov}\left(\Psi(\mathbf{u}^{\prime}),\Psi(\mathbf{u}^{\prime\prime})\right)=\mathbf{k}_{u}(\mathbf{u}^{\prime},\mathbf{u}^{\prime\prime})\in\mathbb{R}^{n_{u}L\times n_{u}L}$, and similarly $\mathrm{cov}\left(\Phi(\mathbf{y}^{\prime}),\Phi(\mathbf{y}^{\prime\prime})\right)=\mathbf{k}_{y}(\mathbf{y}^{\prime},\mathbf{y}^{\prime\prime})\in\mathbb{R}^{n_{y}L\times n_{y}L}$, where $\mathbf{k}_{u}(\cdot,\cdot)$, $\mathbf{k}_{y}(\cdot,\cdot)$ apply $k_{u}(\cdot,\cdot)$, $k_{y}(\cdot,\cdot)$ elementwise, respectively.
@@
$k(\eta^{\prime},\eta^{\prime\prime}):=\left[\Gamma_{1}\ \Gamma_{2}\ -\mathbb{I}\right]\begin{bmatrix}\mathbf{k}_{u}(\mathbf{u}^{\prime},\mathbf{u}^{\prime\prime})&\mathbf{0}\\ \mathbf{0}&\mathbf{k}_{y}(\mathbf{y}^{\prime},\mathbf{y}^{\prime\prime})\end{bmatrix}\begin{bmatrix}\Gamma_{1}^{\top}\\ \Gamma_{2}^{\top}\\ -\mathbb{I}\end{bmatrix}.$ (19)
@@
This completes the Gaussian process model (9) with prior mean and covariance functions (18), (19), noise model (17), and training data (16). To simplify calculation, substituting (18) and (19) into (6)-(8), after some algebraic manipulation, we have
@@
$\mathbf{m}(\boldsymbol{\eta}^{d})$ $=\mathrm{vec}\left(\left[\Gamma_{1}\ \Gamma_{2}\ -\mathbb{I}\right]\begin{bmatrix}\mathbf{m}_{u}(H_{u})\\ \mathbf{m}_{y}(H_{y})\end{bmatrix}\right),$ (20)
@@
$\mathbf{k}(\boldsymbol{\eta}^{d},\eta)$ $=\left(\mathbb{I}\otimes\left[\Gamma_{1}\ \Gamma_{2}\ -\mathbb{I}\right]\right)\mathbf{k}_{uy}(\boldsymbol{\eta}^{d},\eta)\begin{bmatrix}\Gamma_{1}^{\top}\\ \Gamma_{2}^{\top}\\ -\mathbb{I}\end{bmatrix},$ (21)

### v2

@@ 2.1 Data-Driven Prediction for Linear Systems
$\left[\Gamma_{1}\ \Gamma_{2}\right]:=H_{\bar{y}\mathrm{f}}\mathrm{col}\left(H_{\bar{u}},H_{\bar{y}\mathrm{p}}\right)^{\dagger}\in\mathbb{R}^{n_{y}L^{\prime}\times(n_{u}L+n_{y}L_{0})}.$ (3)
{pf}
First note that $(\mathbf{u},\mathbf{y})$ is a trajectory of (1) iff $\left(\Psi(\mathbf{u}),\Phi(\mathbf{y})-\mathbf{e}\right)$ is a trajectory of $G(q)$. According to [26, Proposition 1], when $L_{0}$ is no smaller than the lag of $G(q)$, $\left(\Psi(\mathbf{u}),\Phi(\mathbf{y})-\mathbf{e}\right)$ is a trajectory of $G(q)$ iff $\Phi(\mathbf{y}_{\mathrm{f}})-\mathbf{e}_{\mathrm{f}}=H_{\bar{y}\mathrm{f}}g$, where is $g$ is any solution to
$\begin{bmatrix}H_{\bar{u}}\\ H_{\bar{y}\mathrm{p}}\end{bmatrix}g=\begin{bmatrix}\Psi(\mathbf{u})\\ \Phi(\mathbf{y}_{\mathrm{p}})-\mathbf{e}_{\mathrm{p}}\end{bmatrix}$. (4)
The solution to (4) can be expressed as
$g=\mathrm{col}\left(H_{\bar{u}},H_{\bar{y}\mathrm{p}}\right)^{\dagger}\mathrm{col}\left(\Psi(\mathbf{u}),\Phi(\mathbf{y}_{\mathrm{p}})-\mathbf{e}_{\mathrm{p}}\right)+g_{0},$
where $g_{0}\in\mathrm{null}\left(\mathrm{col}\left(H_{\bar{u}},H_{\bar{y}\mathrm{p}}\right)\right)$. Note that $H_{\bar{y}\mathrm{f}}g_{0}=\mathbf{0}$ from the proof of [13, Theorem 1]. This leads to
$\Phi(\mathbf{y}_{\mathrm{f}})-\mathbf{e}_{\mathrm{f}}=H_{\bar{y}\mathrm{f}}\mathrm{col}\left(H_{\bar{u}},H_{\bar{y}\mathrm{p}}\right)^{\dagger}\mathrm{col}\left(\Psi(\mathbf{u}),\Phi(\mathbf{y}_{\mathrm{p}})-\mathbf{e}_{\mathrm{p}}\right),$
which is equivalent to (2).∎ For the rest of the paper, we define $\bar{\Gamma}_{2}:=\left[\Gamma_{2}\ -\mathbb{I}\right]$ for notational simplicity.
Remark 1.
Equation (2) can be seen as a multi-step-ahead nonlinear ARX model:
$\mathbf{y}_{\mathrm{f}}=\Phi^{-1}\left(\left[\Gamma_{1}\ \Gamma_{2}\right]\mathrm{col}\left(\Psi(\mathbf{u}),\Phi(\mathbf{y}_{\mathrm{p}})\right)-\bar{\Gamma}_{2}\mathbf{e}\right).$ (5)
For linear systems, this predictor is known as the subspace predictor [12], and is equivalent to the solution to the least-squares problem in the noise-free case [13]:
$\underset{\Gamma_{1},\Gamma_{2}}{\mathrm{min}}\ \left\lVert H_{\bar{y}\mathrm{f}}-\left[\Gamma_{1}\ \Gamma_{2}\right]\mathrm{col}\left(H_{\bar{u}},H_{\bar{y}\mathrm{p}}\right)\right\rVert_{F}^{2}.$
If the linear part of the model is known, $\Gamma_{1}$ and $\Gamma_{2}$ can alternatively be derived from state-space parameters [17, Equation 14]:
$\Gamma_{1}=:\left[\Gamma_{11}\ \Gamma_{12}\right],\ \Gamma_{11}=\mathcal{O}_{L^{\prime}}\mathcal{P},\ \Gamma_{12}=\mathcal{T}_{L^{\prime}},\ \Gamma_{2}=\mathcal{O}_{L^{\prime}}\mathcal{Q},$ (6)
where
$\mathcal{C}_{i}$ $:=\begin{bmatrix}A^{i-1}B&\cdots&AB&B\end{bmatrix}$.
This model-based characterization provides insights into hyperprior design later in Section 3.5.
Remark 2.
This work considers an output error structure for $G(q)$ with measurement noise. Alternatively, one can consider an ARX structure with equation errors for $G(q)$, i.e., $\bar{y}_{k}=\beta_{0}^{u}\bar{u}_{k}+\sum_{i=1}^{n_{x}}(\beta_{i}^{u}\bar{u}_{k-i}+\beta_{i}^{y}\bar{y}_{k-i})+e_{k}$. This leads to a different noise model $W^{-1}\mathbf{e}_{\mathrm{f}}$ [9, Proposition 1], where
$W=\begin{bmatrix}1&0&0&\cdots&0\\ -\beta_{1}^{y}&1&0&\cdots&0\\ -\beta_{2}^{y}&-\beta_{1}^{y}&1&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ -\beta_{L^{\prime}-1}^{y}&-\beta_{L^{\prime}-2}^{y}&-\beta_{L^{\prime}-3}^{y}&\cdots&1\\ \end{bmatrix}$,
in place of $\bar{\Gamma}_{2}\mathbf{e}$ in (2). A similar GP regression procedure as in Section 3 can be applied, but with an additional hyperparameter $W$.
2.2 From Basis Functions to Implicit Gaussian Processes
Based on the linear results in the previous subsection, an immediate strategy to analyze H-W systems with unknown input and output nonlinearities is to assume that finite-dimensional basis functions are available for the nonlinearities [5], i.e.,
$\psi(\cdot)$ $=:\sum_{i=1}^{r}a_{i}\psi_{i}(\cdot)=:\left(\mathbf{a}^{\top}\otimes\mathbb{I}\right)\boldsymbol{\psi}(\cdot),$
$\phi(\cdot)$ $=:\sum_{i=1}^{q}b_{i}\phi_{i}(\cdot)=:\left(\mathbf{b}^{\top}\otimes\mathbb{I}\right)\boldsymbol{\phi}(\cdot),$
or equivalently $\psi(\cdot)\in\mathcal{H}_{\psi}$, $\phi(\cdot)\in\mathcal{H}_{\phi}$, where $\mathcal{H}_{\psi}$, $\mathcal{H}_{\phi}$ are the RKHS’s associated to $k_{\psi}(u,u^{\prime}):=\boldsymbol{\psi}(u)^{\top}\boldsymbol{\psi}(u^{\prime})$, $k_{\phi}(y,y^{\prime}):=\boldsymbol{\phi}(y)^{\top}\boldsymbol{\phi}(y^{\prime})$, respectively. Considering the nonlinear system in the lifted input-output space leads to
@@ 2.2 From Basis Functions to Implicit Gaussian Processes
where $\bar{B}:=B\left(\mathbf{a}^{\top}\otimes\mathbb{I}\right)$ and $\bar{D}:=D\left(\mathbf{a}^{\top}\otimes\mathbb{I}\right)$. By considering augmented inputs $\boldsymbol{\psi}(u_{k})$ and outputs $\boldsymbol{\phi}(y_{k})$, (7) admits linearity in the sense that for any two input-output trajectories of (7), their linear combinations are still valid trajectories. However, as also pointed out in [5], (7) is not a standard linear model since the operator $\left(\mathbf{b}^{\top}\otimes\mathbb{I}\right)\in\mathbb{R}^{n_{y}\times n_{y}q}$ is not invertible. This prohibits the use of WFL to quantify all possible input-output trajectories and provide data-driven predictions when unknown nonlinear output uncertainties exist. Kernelized reformulations of WFL are presented in [29, 23]. However, similar issues restrict its application to Wiener systems.
Even when we restrict ourselves to Hammerstein systems, a finite-dimensional basis function decomposition is often difficult to obtain, and the application of WFL requires persistency of excitation in the lifted signal space, which becomes difficult as the basis function dimension increases [4]. Although an infinite-dimensional function space associated to a general kernel can be applied with the kernelized reformulation, the persistency of excitation condition can no longer be satisfied, and exact trajectory certification becomes impossible [16].
Since a straightforward application of WFL is difficult, an alternative strategy is to directly learn the nonlinear ARX model (5) for H-W systems. Yet, to the best of our knowledge, existing function learning methods cannot encode the composite nonlinear structure in (5). One can, of course, disregard the specific structure in (5) and learn a purely black-box model $\mathbf{y}_{\mathrm{f}}=f(\mathbf{u},\mathbf{y}_{\mathrm{p}})$ by e.g., basis function decomposition [22], kernel regression [16], or GP regression [19]. However, this would lead to deteriorated performance since a much larger function space is learned from, and the predictor would not correspond to an H-W system.
Due to the above reasons, in this work, we employ the implicit trajectory characterization (2) instead of the explicit one (5), since (2) does not contain any composite of nonlinear functions. This enables a physics-informed design that incorporates the H-W structure in learning. Implicit GP regression is applied to jointly learn the linear mappings $\Gamma_{1}$, $\Gamma_{2}$ and the nonlinearities $\psi(\cdot)$, $\phi(\cdot)$, as will be detailed in the following section.
3 Implicit Gaussian Process Data-Driven Prediction
This section presents the implicit GP data-driven prediction scheme for H-W systems. Section 3.1 starts by explaining the standard procedure of GP regression, where the structured kernel design for the implicit GP is obtained in Section 3.2. Section 3.3 incorporates monotonicity information in the regression by expectation propagation. The posterior predictions from the implicit model are given in Section 3.4. Section 3.5 investigates the hyperparameter tuning problem for the linear model parameters, followed by Section 3.6, which presents a scheme to recover the nonlinear functions.
3.1 Gaussian Process Regression
@@
The standard procedure of GP regression is briefly summarized as follows [35]. Consider the problem of learning an unknown function $f:\mathbb{R}^{n_{\eta}}\rightarrow\mathbb{R}^{n_{\chi}}$ from data $\left(\boldsymbol{\boldsymbol{\eta}}^{d},\boldsymbol{\chi}^{d}\right)=\left(\eta_{k}^{d},\chi_{k}^{d}\right)_{k=1}^{M}$, where $\chi^{d}_{k}=f(\eta^{d}_{k})+\epsilon_{k}$ and $\epsilon_{k}$ is zero-mean i.i.d. Gaussian with covariance $\Sigma$. The function $f$ is modeled as a GP with prior mean $m(\cdot)$, prior covariance $k(\cdot,\cdot;\Theta)$, and hyperparameters $\Theta$, i.e., $f(\cdot)\sim\mathcal{GP}\left(m(\cdot),k(\cdot,\cdot;\Theta)\right)$. To predict $f(\eta)$ at any query point $\eta$, the joint distribution of $f(\eta)$ and $\boldsymbol{\chi}^{d}$ is given by the following Gaussian distribution:
[…]
@@
$f(\eta)|\boldsymbol{\chi}^{d}\sim\mathcal{N}\left(m_{\mathrm{p}}(\eta),k_{\mathrm{p}}(\eta)\right),$ (11)
@@
$m_{\mathrm{p}}(\eta)$ $:=m(\eta)+\mathbf{k}^{\top}(\boldsymbol{\eta}^{d},\eta)\left(K+\bar{\Sigma}\right)^{-1}\left(\boldsymbol{\chi}^{d}-\mathbf{m}(\boldsymbol{\eta}^{d})\right),$ (12)
$k_{\mathrm{p}}(\eta)$ $:=k(\eta,\eta)-\mathbf{k}^{\top}(\boldsymbol{\eta}^{d},\eta)\left(K+\bar{\Sigma}\right)^{-1}\mathbf{k}(\boldsymbol{\eta}^{d},\eta),$ (13)
where the dependence of $K$ on $\boldsymbol{\eta}^{d}$ is omitted for brevity. Then, $f(\eta)$ is predicted as the posterior mean $m_{\mathrm{p}}(\eta)$.
@@
$\Theta$ $=\mathrm{arg}\underset{\Theta}{\mathrm{max}}\ p\left(\boldsymbol{\chi}^{d}|\Theta\right)$ (14)
$=\mathrm{arg}\underset{\Theta}{\mathrm{min}}\ \mathrm{logdet}\left(\Lambda(\Theta)\right)+\delta^{\top}\Lambda^{-1}(\Theta)\delta,$ (15)
where $\delta:=\boldsymbol{\chi}^{d}-\mathbf{m}(\boldsymbol{\eta}^{d})$ and $\Lambda(\Theta):=K(\boldsymbol{\eta}^{d},\boldsymbol{\eta}^{d};\Theta)+\bar{\Sigma}$.
3.2 Structured Kernel Design
Since the input and output nonlinearities are unknown, they are modeled as GPs with prior distributions:
Remark 3.
Most standard kernel structures include a scaling parameter as a hyperparameter. In our setting, the model (1) is ambiguous up to scaling factors, i.e., the model triplet $\left(\psi(\cdot),G(q),\phi^{-1}(\cdot)\right)$ is equivalent to $\left(a\psi(\cdot),\frac{1}{ab}G(q),b\phi^{-1}(\cdot)\right)$, $a,b\in\mathbb{R}$. Thus, the scaling parameter can be fixed to 1 without loss of generality. Similarly, the mean functions $m_{u}$ and $m_{y}$ can also be normalized.
Remark 4.
Although the prior distribution of $\Gamma_{1},\Gamma_{2}$ can be derived from (3) and the Gaussian prior (16), it is dependent on $\psi(\cdot)$ and $\phi(\cdot)$ in a nonlinear way, and thus the Gaussianity is not preserved. Instead, we opt for directly learning $\Gamma_{1},\Gamma_{2}$ in a hierarchical model by regarding them as hyperparameters, as will be discussed in Section 3.5.
The implicit GP regression problem is formulated as follows. Consider the trajectory characterization equation (2). Let $\eta=\mathrm{col}\left(\mathbf{u},\mathbf{y}\right)$ and
$f(\eta)=\left[\Gamma_{1}\ \bar{\Gamma}_{2}\right]\mathrm{col}\left(\Psi(\mathbf{u}),\Phi(\mathbf{y}_{\mathrm{p}}),\Phi(\mathbf{y}_{\mathrm{f}})\right),$ (17)
where $n_{\eta}=(n_{u}+n_{y})L$ and $n_{\chi}=n_{y}L^{\prime}$. We would like to train a GP model of $f(\eta)$ by using data $\left(\boldsymbol{\boldsymbol{\eta}}^{d},\boldsymbol{\chi}^{d}\right)$ where M=NL+1M=N-L+1,
@@ 3.2 Structured Kernel Design
$\bar{\Sigma}$ $=\mathbb{I}\otimes\mathrm{cov}\left(\bar{\Gamma}_{2}\mathbf{e}\right)=\sigma^{2}\left(\mathbb{I}\otimes\left(\Gamma_{2}\Gamma_{2}^{\top}+\mathbb{I}\right)\right).$ (19)
This is known as an implicit GP model [27].
The prior mean and covariance functions $m(\cdot)$ and $k(\cdot,\cdot)$ for this implicit GP model are then derived from the GP characterization of $\psi(\cdot)$ and $\phi(\cdot)$ in (16). The mean function of an input-output trajectory $\eta=\mathrm{col}\left(\mathbf{u},\mathbf{y}\right)$ is given by
$m(\eta)=\left[\Gamma_{1}\ \bar{\Gamma}_{2}\right]\mathrm{col}\left(\mathbf{m}_{u}(\mathbf{u}),\mathbf{m}_{y}(\mathbf{y})\right),$ (20)
where $\mathbf{m}_{u}(\cdot)$, $\mathbf{m}_{y}(\cdot)$ apply $m_{u}(\cdot)$, $m_{y}(\cdot)$ elementwise, respectively. Consider two input-output trajectories $\eta^{\prime}=\mathrm{col}\left(\mathbf{u}^{\prime},\mathbf{y}^{\prime}\right)$ and $\eta^{\prime\prime}=\mathrm{col}\left(\mathbf{u}^{\prime\prime},\mathbf{y}^{\prime\prime}\right)$. The covariance matrix between $\Psi(\mathbf{u}^{\prime})$ and $\Psi(\mathbf{u}^{\prime\prime})$ is given by $\mathrm{cov}\left(\Psi(\mathbf{u}^{\prime}),\Psi(\mathbf{u}^{\prime\prime})\right)=\mathbf{k}_{u}(\mathbf{u}^{\prime},\mathbf{u}^{\prime\prime})\in\mathbb{R}^{n_{u}L\times n_{u}L}$, and similarly $\mathrm{cov}\left(\Phi(\mathbf{y}^{\prime}),\Phi(\mathbf{y}^{\prime\prime})\right)=\mathbf{k}_{y}(\mathbf{y}^{\prime},\mathbf{y}^{\prime\prime})\in\mathbb{R}^{n_{y}L\times n_{y}L}$, where $\mathbf{k}_{u}(\cdot,\cdot)$, $\mathbf{k}_{y}(\cdot,\cdot)$ apply $k_{u}(\cdot,\cdot)$, $k_{y}(\cdot,\cdot)$ elementwise, respectively.
@@
$k(\eta^{\prime},\eta^{\prime\prime}):=\Gamma_{1}\mathbf{k}_{u}(\mathbf{u}^{\prime},\mathbf{u}^{\prime\prime})\Gamma_{1}^{\top}+\bar{\Gamma}_{2}\mathbf{k}_{y}(\mathbf{y}^{\prime},\mathbf{y}^{\prime\prime})\bar{\Gamma}_{2}^{\top}.$ (21)
@@
To simplify calculation, substituting (20) and (21) into (8)-(10), after some algebraic manipulation, we have
[…]
@@
$\mathbf{k}(\boldsymbol{\eta}^{d},\eta)$ $=\left(\mathbb{I}\otimes\Gamma_{1}\right)\mathbf{k}_{u}(\mathrm{vec}\left(H_{u}\right),\mathbf{u})\Gamma_{1}^{\top}$
@@
$\quad+\left(\mathbb{I}\otimes\bar{\Gamma}_{2}\right)\mathbf{k}_{y}(\mathrm{vec}\left(H_{y}\right),\mathbf{y})\bar{\Gamma}_{2}^{\top},$ (23)